OpenStudy (anonymous):

The area of a rectangular piece of land is 240 square meters. If the length of the land was 5 meters less and the width was 3 meters more, the shape of the land would be a square. Part A: Write an equation to find the width (x) of the land. Show the steps of your work. (5 points) Part B: What is the width of the land in meters? Show the steps of your work. (5 points)

OpenStudy (anonymous):

@sidsiddhartha @AliyaMarie9 @arbershabani97

OpenStudy (campbell_st):

well if the length is l and width is w, and side length of the square is x then for length of the square x = l - 5 or making l the subject l = x + 5 same thinking for the width x = w + 3 or by making w the subject w = x - 3 And you know the area is 240 = lw which means the equation of the square can be written as 240 = (x + 5)(x - 3) so now you have the equation... all you need to do is solve for x. by forming the quadratic equation you can solve.

OpenStudy (anonymous):

x=-15,17?

OpenStudy (anonymous):

oh wait switch the signs

OpenStudy (anonymous):

x=15,-17

OpenStudy (campbell_st):

thats better... so which solution would you use...?

OpenStudy (anonymous):

do i subtract 5 and add 3

OpenStudy (anonymous):

i don't really get what ur asking

OpenStudy (campbell_st):

well the question is poorly worded... but I'd assume you need to go to the original width equation and substitute the appropriate value of x.

OpenStudy (anonymous):

so 5 less then 15 is 10 and 3 more then -17 is -14 but idk i don't get it

OpenStudy (campbell_st):

well let me ask you this... you are dealing with measurements.... can you have negative measurements when calculating area.... it may help you decide which solution to use....

OpenStudy (campbell_st):

and remember for the original area length = x + 5 width = x - 3

OpenStudy (anonymous):

okay so it would definitely be 15 15+5 is 20 15-3 is 12 12 times 20 is 240 and wallah

OpenStudy (anonymous):

so that's all for part a or for both?

OpenStudy (anonymous):

the width is 12 meters right?

OpenStudy (campbell_st):

that looks good for me...